OpenStudy (durply):

which restriction is appropriate for the variable 3y^2-3y/y+1

OpenStudy (durply):

@sammixboo

OpenStudy (legomyego180):

what can you not divide by in this equation, or in any equation

OpenStudy (campbell_st):

well it depends on the restrictions you have been taught about are if you factor the numerator \[\frac{3y^2 - 3y}{y + 1} = \frac{3(y -1)(y + 1)}{(y + 1)}\] the the graph doesn't have a vertical asymptote which is a discontinuity expression only had a point of discontinuity... and discontinuities occur when the denominator is zero... so find the value of y that makes the denominator zero

OpenStudy (agent0smith):

@campbell_st it's just an expression, not a function, so it's only asking about the restriction for the denominator.